# Near-Field Localization and Phase Shift Optimization for RIS-Assisted Non-Ideal OFDM Systems

### v2

@@ Abstract
By incorporating reconfigurable intelligent surface (RIS) into communication-assisted localization systems, the issue of signal blockage caused by obstacles can be addressed, and passive beamforming can be employed to enhance localization accuracy. However, existing works mainly consider ideal channels and do not account for the effects of realistic impairments like carrier frequency offset (CFO) and phase noise (PN) on localization. This paper proposes an iterative joint estimation algorithm for CFO, PN, and user position based on maximum a posteriori (MAP) criterion and gradient descent (GD) algorithm. Closed-form expressions for CFO and PN updates are provided. The hybrid Cramér-Rao lower bound (HCRLB) for the estimation parameters is derived, and the ambiguity in CFO and PN estimation is analyzed. To minimize the HCRLB, a non-convex RIS shift optimization problem is formulated and is transformed into a convex semidefinite programming (SDP) problem using the technique of semidefinite relaxation (SDR) and Schur complement. After optimizing the RIS phase shift, the theoretical positioning accuracy within the area of interest (AOI) can be improved by two orders of magnitude, with a maximum positioning root mean square error (RMSE) lower than $\rm 10^{-2}m$.
@@
Near-field localization, reconfigurable intelligent surface, passive beamforming, carrier frequency offset, phase noise.
@@ I Introduction
With the widespread application of fifth-generation (5G) networks worldwide, research focus has shifted towards future six-th generation (6G) mobile communication technologies. For 5G mobile networks, the demand for positioning accuracy has already reached sub-meter levels due to the need to support various location-based services. In 6G networks, it is expected that high-precision positioning at the centimeter level will play an even more important role [1, 2, 3]. While global navigation satellite systems (GNSSs) have been widely used in various positioning scenarios [25], they suffer from issues such as insufficient positioning accuracy and signal instability due to obstacles [26, 27]. In recent years, research has been conducted on using communication signals to assist in positioning, aiming to achieve more accurate and continuous positioning alongside GNSS signals. For example, Shahmansoori et al. [10] proposed a three-stage positioning algorithm that improves the simultaneous orthogonal matching pursuit (SOMP) algorithm and the space alternating generalized expectation-maximization (SAGE) algorithm, and exploited the sparsity of millimeter-wave multiple-input multiple-output (MIMO) channels to lower the complexity of the channel model. Hu et al. [11] presented a two-dimensional (2D) positioning algorithm based on the estimating signal parameters via rotational invariance technique (ESPRIT), which estimates user positions by constructing signal subspaces. The estimator has closed-form expressions, which reduces the computational complexity significantly compared to traditional estimation methods. However, in these methods that utilize communication signals for localization, the problem of signal blockage by obstacles still remains unavoidable.
In recent years, reconfigurable intelligent surface (RIS) technology has emerged as a key enabling technology for 6G networks [4, 5, 23]. This technology allows for reconfiguring wireless channel environments, enabling communication even when the line-of-sight (LoS) link is obstructed. Introducing the RIS to assist in communication-based positioning has become a research focus [32, 31]. From the perspective of mobility, the research can be divided into two categories: static positioning [17, 18, 19] and dynamic positioning [15, 13]. The former considers scenarios where the user undergoes minimal movement within a coherent time, and the latter involves scenarios with the opposite characteristic. From the perspective of positioning distance, the research can also be divided into two categories: far-field positioning [19, 13, 20] and near-field positioning [18, 19, 13, 24]. In the far-field case, the distance between the user and the RIS exceeds Fraunhofer distance [8], and the signal can be approximated as a planar wave. Conversely, in the near-field case, the signal cannot be treated as parallel propagation and should be considered a spherical wave.
@@
For example, Luan et al. [12] focused on the near-field static 2D positioning scenario. The authors proposed a robust optimization method to optimize the RIS phase shift for improving localization accuracy. The paper also investigated the power optimization problem of the localization system. Elzanaty et al. [13] studied RIS-assisted localization for multi-antenna mobile users. They proposed a three-dimensional (3D) model that can be applied in both far-field and near-field scenarios and analyzed the position and orientation error bounds. Lin et al. [14] introduced a novel twin-RIS configuration comprising two RISs with relative spatial rotation. To extract the parameters of the cascaded channel, the authors employed array signal processing and atomic norm denoising techniques. They established nonlinear equations based on the characteristics of the twin-RIS structure and accurately recovered the multipath parameters. User localization for static users in the far-field scenario is then performed based on the estimated channel parameters. Teng et al. [15] introduced a new message-passing algorithm called BULT to effectively solve the localization problem for mobile users in the far-field scenario. They also derived the Bayesian Cramér-Rao lower bound for the proposed algorithm.
@@
Although there has been extensive research on user localization with RIS in various scenarios such as static, dynamic, far-field, and near-field, most of the existing works assumed ideal channel conditions without considering realistic impairments like carrier frequency offset (CFO) and phase noise (PN) that can introduce uncertainties and affect system performance. Only a few studies have considered modeling certain interferences. For example, [13, 20, 16] took into account clock offset, and [16] considered phase offset. However, to the best of our knowledge, there is currently no research that specifically analyzed CFO and PN in RIS-assisted orthogonal frequency division multiplexing (OFDM) localization systems.
In an OFDM system, CFO and PN can convolve with the data symbols, causing signal constellation rotation and introducing inter-carrier interference (ICI). This interference significantly degrades the overall system performance. Moreover, most aforementioned works have assumed the subcarriers’ signals are independent, which cannot hold in the presence of CFO and PN due to the absence of ICI. Therefore, localization algorithms need to leverage the received time-domain signals to account for these effects and accurately estimate the user’s position in this case.
In this paper, we consider the problem of joint estimation of CFO, PN, and user position in a RIS-assisted OFDM system. The contributions of this paper are outlined as follows.
  • We investigate the problem of static near-field localization in an OFDM system assisted by RIS, considering the effects of CFO and PN. Moreover, we establish a time-domain model and introduce the frequency-domain channel with time-delay parameters into the model through time-frequency transformation to meet the requirements of position estimation.
  • We propose an iterative estimation approach based on the maximum a posteriori (MAP) criterion and gradient descent (GD) algorithm to jointly estimate the CFO, the PN, and the user position. The algorithm significantly reduces complexity by deriving closed-form expressions for CFO and PN estimation.
@@
  • We derive the hybrid Cramér-Rao lower bound (HCRLB) for the joint estimation problem and analyze the phase ambiguity issue in CFO and PN estimation. A new metric is introduced to evaluate the accuracy of joint CFO and PN estimation.
  • We implement phase shift optimization for RIS to enhance the theoretical positioning accuracy. By applying the semidefinite relaxation (SDR) method and utilizing the Schur complement, we convert the initially non-convex optimization problem into a convex semidefinite programming (SDP) problem. It is shown that the theoretical positioning accuracy has been improved by two orders of magnitude.
Figure 1: RIS-aided 3D localization scenario, where the LoS link is blocked.
@@
The paper is structured as follows. Section II introduces the system model, while Section III presents the joint estimation algorithm. Section IV is dedicated to the derivation of the HCRLB for the proposed joint estimation problem. The proposed RIS phase shift optimization method is discussed in Section V. Section VI showcases comprehensive simulation results, and the paper concludes in Section VII.
@@
Notations: $\mathbb{E}(\cdot)$, $p(\cdot)$ denote the expectation and the probability density function (PDF) of the argument, respectively. ${\rm tr}(\mathbf{M})$, $\det\mathbf{M}$, $\mathbf{M}^{\top}$, $\mathbf{M}^{*}$, $\mathbf{M}^{\rm H}$, $\mathbf{M}^{\rm-1}$ denote the trace, determinant, transpose, conjugate, hermitian, and inverse of matrix $\mathbf{M}$, respectively. ${\rm Diag}(\bm{v})$ indicates a diagonal matrix created by assigning the elements of vector $\bm{v}$ along its main diagonal. $\bm{v}[n]$ denotes the $n$-th entry of vector $\bm{v}$. Unless otherwise specified, all indices in the article start from 0. ${\rm Blkdiag}(\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1})$ denotes a block diagonal matrix with $\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1}$ being its diagonal matrices. $\left[\mathbf{M}\right]_{i:j,:}$ refers to a submatrix encompassing rows $i$ to $j$ and all columns of matrix $\mathbf{M}$, while $\left[\mathbf{M}\right]_{:,i:j}$ denotes a submatrix encompassing columns $i$ to $j$ and all rows of matrix $\mathbf{M}$. $\bm{0}_{M\times N}$, and $\mathbf{I}_{M}$ denote $M\times N$ zeros matrix, and $M\times M$ identity matrix, respectively. ${}\Re\left\{\cdot\right\}$ and ${}\Im\left\{\cdot\right\}$ represent the real and imaginary components of the complex value, respectively. $\mathbb{C}^{m\times n}$ and $\mathbb{R}^{m\times n}$ denote $m\times n$ matrices with complex and real entries, respectively. Finally, $\mathcal{CN}(\bm{\mu},\mathbf{\Sigma})$ and $\mathcal{N}(\bm{\mu},\mathbf{\Sigma})$ represent complex and real Gaussian distributions, respectively, characterized by a mean vector $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$.
@@ II System Model
Fig. 1 illustrates the RIS-assisted OFDM downlink wireless communication system under consideration, comprising a single anchor node (AN) and a single user equipment (UE). In this scenario, we make the assumption that there are obstacles blocking the LoS path between AN and UE, and the channel state information (CSI) remains unchanged during the localization process, i.e., within the duration of one OFDM symbol, similar to [19, 12, 16]. The positions of the AN and each element of the RIS are assumed to be known, which can be denoted as $\bm{p}_{a}$, and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, where $r$ represents the index of the RIS element among a total of $N_{R}$ elements. The position of UE, which is the primary parameter to be estimated, is denoted as $\bm{p}_{u}=[x_{u},y_{u},z_{u}]^{\top}$.
The RIS elements are organized in the form of a uniform planar array (UPA). We take the bottom-left corner of the RIS as the reference point and place it at the coordinate origin. The RIS is positioned on the $yz$ plane, and both edges on the coordinate axes are located in the positive half-axis. The coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times((r-1)\mod N_{R}),d\times\lceil\frac{r}{\sqrt{N_{R}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of RIS, whose value is typically set as half of the wavelength. To satisfy the conditions of the near-field setup, both the distance between AN and RIS, as well as the distance between RIS and UE, should lie within the Fresnel region [8]. Specifically, the distances should lie within the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{N_{R}}-1)$ represents the aperture of RIS and $\lambda$ is the operating wavelength. Additionally, we adopt a polar coordinate system for estimating the user’s position, i.e., using $d_{Ru}$, $\phi_{az}$, $\phi_{el}$ to express the user’s position, which represents the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively. The transformation relationships between these three parameters and $x_{u}$, $y_{u}$, $z_{u}$ are as follows
$d_{Ru}$ $=\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}},$ (1a)
$\phi_{az}$ $=\arctan\frac{y_{u}}{x_{u}},$ (1b)
@@
The localization process in the paper is performed using one OFDM symbol, with $N$ subcarriers, a center carrier frequency of $f_{c}$, corresponding to a wavelength of $\lambda$, and a signal bandwidth of $B$. Upon sampling at a rate of $T_{s}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at UE, the received time-domain signal can be represented as
$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\tilde{\bm{h}}+\tilde{\bm{v}},$ (3)
@@
  • ${\bm{y}\triangleq\left[\bm{y}[0],\bm{y}[1],\cdots,\bm{y}[N-1]\right]}^{\top}$,
  • $P$ is the transmit power,
  • $\bm{\Lambda}_{\phi}$ is the impact caused by CFO on the signal. It is a diagonal matrix, and each item on the diagonal can be expressed as $\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n\Delta f}{NB}}=e^{j2\pi\frac{n}{N}\phi}$, where $\phi=\frac{\Delta f}{B}$ is called the normalized CFO,
  • $\bm{\Lambda}_{\bm{\theta}}\triangleq{\rm Diag}\left[e^{j\bm{\theta}[0]},e^{j\bm{\theta}[1]},\cdots,e^{j\bm{\theta}[N-1]}\right]$ is the impact caused by PN on the signal, further details about which will be discussed in the later part of this section,
  • $\mathbf{S}\triangleq\sqrt{N}\mathbf{F}^{\rm H}{\rm Diag}(\bm{s})$, where $\mathbf{F}^{\rm H}$ is the inverse fast Fourier transform (IFFT) matrix and $\bm{s}$ is the frequency domain transmitted symbol sequence at the AN,
  • $\tilde{\bm{h}}\triangleq\bm{h}_{LoS}+\bm{h}_{NLoS}$ is the frequency domain channel, with $\bm{h}_{LoS}$ and $\bm{h}_{NLoS}$ representing the LoS and the non-line-of-sight (NLoS) part of the channel, respectively. We assume $\bm{h}_{NLoS}$ follows a known distribution $\bm{h}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{I}_{N})$. Hence, The impact caused by $\bm{h}_{NLoS}$ on the signal can be combined with the additive Gaussian noise $\tilde{\bm{v}}$. Consequently, we can rewrite (3) as

### v3

# Reconfigurable Intelligent Surface-Assisted Localization in OFDM Systems with Carrier Frequency Offset and Phase Noise

@@ Abstract
Reconfigurable intelligent surface (RIS)-assisted communication systems have been extensively studied for providing high-precision location services. However, most studies have overlooked the impact of carrier frequency offset (CFO) and phase noise (PN) resulting from hardware impairments on localization. This paper presents a novel, alternating optimization (AO)-based algorithm to jointly estimate the CFO, PN, and user position, where, provided the user position, closed-form expressions for the CFO and PN are derived per iteration, significantly reducing the complexity and enhancing the stability of the algorithm. Another important aspect is a new RIS phase shift optimization algorithm developed to minimize the analytical lower bound of localization accuracy, hence benefiting localization. The semidefinite relaxation method and Schur complement are utilized to convexify this challenging non-convex optimization problem to a semidefinite program. Simulations demonstrate the effectiveness of the proposed algorithms, with the localization accuracy enhanced by two orders of magnitude. The localization accuracy of the proposed algorithm is close to the analytical lower bound, with a root mean square error of lower than $\rm 10^{-2}\>m$.
@@
Localization, reconfigurable intelligent surface, phase shift configuration, carrier frequency offset, phase noise.
@@ I Introduction
I-A Motivation and Challenges
With the development of industrial Internet of Things, there is an increasing demand for centimeter-level localization accuracy in the sixth-generation (6G) mobile networks [1, 2, 3]. While global navigation satellite systems (GNSSs) have been widely applied to outdoor scenarios [4], they undergo insufficient localization accuracy and signal instability arising from obstacles [5, 6]. In recent years, communication systems have been considered to assist in localization, e.g., integrated sensing and communication (ISAC), aiming to achieve more accurate and continuous localization alongside GNSS signals.
Although communication-assisted localization systems have been extensively studied, few studies have considered the adverse effects of hardware impairments, such as carrier frequency offset (CFO) and phase noise (PN), on localization accuracy. CFO is a constant parameter due to the frequency offset between the transmitter and the receiver. Doppler shift can also lead to CFO. PN is a random process resulting from the fluctuation of the oscillators at the transmitters and receivers [7]. CFO and PN can result in signal constellation rotations in single carrier communication systems, or inter-carrier interference (ICI) in broadband communication systems, for example, orthogonal frequency division multiplexing (OFDM) systems. The overall communication performance and the localization accuracy can degrade significantly, due to the CFO and PN.
Another challenge arises from the increasingly likely blockages of signals, e.g., millimeter wave (mmWave) signals with the quasi-optical property, in many dense urban areas [8]. Reconfigurable intelligent surface (RIS) [9, 10] has been considered in communication systems to deal with the problem of signal blockage. Localization can be implemented with the help of virtual line-of-sight (LoS) paths provided by RISs even when the LoS paths are blocked, and can be improved via passive beamforming. In this sense, RIS-assisted user localization is of practical interest, and could be particularly challenging with the further consideration of the CFO and PN. However, this has not been studied in the existing literature.
I-B Related works
Communication systems have been widely employed to assist in high-precision user localization, typically under a far-field setting; i.e., the user and the anchors are in the far field of each other. For example, Shahmansoori et al. [11] proposed a three-stage localization algorithm that improves the simultaneous orthogonal matching pursuit algorithm and the space alternating generalized expectation-maximization algorithm, and exploited the sparsity of millimeter-wave multiple-input multiple-output (MIMO) channels to simplify the channel model. Hu et al. [12] presented a two-dimensional (2D) localization algorithm based on the celebrated estimating signal parameters via rotational invariance technique (ESPRIT), which estimates user positions by constructing signal subspaces.
Some studies [13, 14] have explored the impact of realistic hardware impairments, including the CFO and PN, with no RIS considered. For instance, Meles et al. [13] simulated the effects of CFO and PN on angle-of-arrival (AOA) estimation. Chen et al. [14] considered the impacts of CFO, PN, mutual coupling, and power amplifier nonlinearity on 6G localization, and performed analysis using the derived misspecified Cramér-Rao lower bound (CRLB). However, neither of them provided an estimation algorithm to handle the hardware impairments.
Some other works [15, 16, 17, 18] have delved into localization algorithm design for scenarios with either CFO or PN. On the one hand, the CFO was studied in [15, 16]. Qiu et al. [15] designed an maximum likelihood estimation (MLE)-based weighted iterative algorithm for the joint estimation of CFO and AOA, yet the coupling of CFO and position parameters in each iteration incurs high complexity. Zakariyya et al. [16] proposed a high-resolution multiple signal classification (MUSIC) algorithm for localization, but did not utilize the CFO estimates in localization, leading to a loss in localization accuracy. On the other hand, the PN was considered in [17, 18]. Xiao et al. [17] combined the expectation maximization (EM) and MUSIC algorithms for joint estimation of AOA and PN, but only considered a single-carrier system under a 2D setting. Zheng et al. [18] formulated a maximum a-posteriori (MAP) problem iteratively solved using the block majorization minimization (MM) algorithm after approximation, where the PN was simplified as a Gaussian variable. However, the above methods have not considered the use of RISs nor addressed the signal blockage by obstacles.
RIS has been integrated into localization systems to address the problem of signal blockage, but little consideration has been given to the above-mentioned hardware impairments, i.e., CFO and PN. From the perspective of user mobility, existing studies can be divided between static localization [19, 20, 21] and dynamic localization [22, 23]. The former encompasses scenarios where the users undergo minimal movement within a coherent time. The latter involves scenarios when the users move at a relatively high speed, and the Doppler shift cannot be overlooked. From the perspective of localization distance, existing research can be divided between far-field localization [21, 23, 24] and near-field localization [20, 21, 23, 25]. In the far-field case, the user-RIS distance exceeds Fraunhofer distance [26], and the signal can be approximated as a planar wave. Conversely, in the near-field case, the signal cannot be treated as parallel propagation and should be considered as a spherical wave.
@@ I-B Related works
For example, Luan et al. [27] focused on the near-field static 2D localization scenario, and proposed a robust optimization algorithm to configure the RIS phase shifts for improving localization accuracy. The transmit powers were also optimized in the considered localization system. Ozturk et al. [28] further introduced a more practical RIS model with phase-dependent amplitude variations in the near-field circumstance. The misspecified CRLB was derived and an approximated mismatched maximum likelihood-based algorithm was proposed for joint UE localization and RIS parameter estimation. The system model in this paper also belongs to the near-field static category. Although we adopt the classical ideal RIS phase shift model with unit amplitude for simplicity, we first take the CFO and PN into consideration. Elzanaty et al. [23] studied RIS-assisted localization for multi-antenna mobile users. They proposed a three-dimensional (3D) model that applies to both far-field and near-field scenarios, and analyzed the position and orientation error bounds. Lin et al. [29] introduced a novel twin-RIS configuration comprising two RISs with relative spatial rotations. To extract the parameters of the cascaded channel, the authors employed array signal processing and atomic norm denoising techniques. They established nonlinear equations exploiting the characteristics of the twin-RIS structure to recover the multipath parameters. User localization for static users in the far-field scenario was then performed based on the estimated channel parameters. Teng et al. [22] introduced a new message-passing algorithm called BULT to locate mobile users in the far-field scenario. The Bayesian CRLB was derived for BULT.
@@
The existing RIS-assisted localization studies have assumed ideal channel conditions without considering realistic hardware impairments. Only several studies have considered modeling some specific impairments. For example, the studies in [23, 24, 30] took into account clock offset, and the work in [30] considered phase offsets. Zhang et al. [31] considered the CFO while overlooking the effects of PN. They proposed a MUSIC-based algorithm that directly estimates the position parameters together with the CFO. However, the algorithm is only suitable for far-field scenarios. To the best of our knowledge, no existing research has specifically analyzed the impact of both CFO and PN on RIS-assisted near-field localization systems.
I-C Contributions
In this paper, we investigate a RIS-assisted OFDM system, where practical CFO and PN convolve with the data symbols, causing signal constellation rotation and ICI and, consequently, hindering user localization in near-field environments. The key contributions of this paper are outlined as follows:
  • We investigate a problem of static near-field localization in a RIS-aided OFDM system, while taking into account the CFO and PN. We propose an alternating optimization (AO)-based iterative estimation approach, where the MAP criterion and gradient descent (GD) algorithm are employed with closed-form expressions derived for the CFO and PN estimation.
@@ I-C Contributions
  • We derive the hybrid CRLB (HCRLB) for the considered joint estimation problem. The phase ambiguity in CFO and PN estimation is analyzed. To evaluate the accuracy of joint CFO and PN estimation, a joint mean square error (MSE) metric is introduced.
  • We further optimize the phase shifts of the RIS to improve the performance of localization. By applying the semi-definite relaxation (SDR) and the Schur complement, we convert the non-convex optimization problem into a convex semi-definite programming (SDP) problem solved efficiently.
  • We perform comprehensive numerical simulations to assess the performance of the proposed algorithms. It is shown that the localization accuracy is improved by two orders of magnitude compared to random RIS phase shift configuration. Compared to the state-of-the-art benchmarks, our algorithms can improve the localization accuracy by about one order of magnitude under relatively low transmit power and even by two orders of magnitude under high transmit power.
Figure 1: The considered RIS-aided 3D near-field localization system, including an AN, a UE, and an $N_{R}$-element RIS. The LoS link between the AN and UE is obstructed. The UE utilizes the signals transmitted by the AN and reflected by the RIS to localize its position.
@@
The rest of this paper is organized as follows. Section II presents the system model. Section III presents the joint estimation algorithm. Section IV derives the HCRLB for the joint estimation algorithm. The proposed RIS phase shift optimization method is delineated in Section V. Section VI showcases comprehensive simulation results, and the paper is concluded in Section VII.
@@
Notation: $\mathbb{E}(\cdot)$ and $p(\cdot)$ denote expectation and probability density function (PDF), respectively. ${\rm tr}(\cdot)$, $\det(\cdot)$, $(\cdot)^{\top}$, $(\cdot)^{*}$, $(\cdot)^{\rm H}$, and $(\cdot)^{\rm-1}$ denote trace, determinant, transpose, conjugate, hermitian, and inverse, respectively. ${\rm Diag}(\bm{v})$ indicates a diagonal matrix where the entries of vector $\bm{v}$ appear on its main diagonal. $\bm{v}[n]$ denotes the $n$-th entry of vector $\bm{v}$. ${\rm Blkdiag}(\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1})$ denotes a block diagonal matrix with $\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1}$ being the matrices along its main diagonal. $\left[\mathbf{M}\right]_{i:j,:}$ refers to a sub-matrix encompassing rows $i$ to $j$ of matrix $\mathbf{M}$, while $\left[\mathbf{M}\right]_{:,i:j}$ denotes a sub-matrix encompassing columns $i$ to $j$ of matrix $\mathbf{M}$. $\bm{0}_{M\times N}$ and $\mathbf{I}_{M}$ denote the $M\times N$ all-zero matrix and the $M\times M$ identity matrix, respectively. ${}\Re\left\{\cdot\right\}$ and ${}\Im\left\{\cdot\right\}$ represent the real and imaginary components of a complex value, respectively. $\mathbb{C}^{m\times n}$ and $\mathbb{R}^{m\times n}$ collect all $m\times n$ matrices with complex and real entries, respectively. $\mathcal{CN}(\bm{\mu},\mathbf{\Sigma})$ and $\mathcal{N}(\bm{\mu},\mathbf{\Sigma})$ represent complex and real Gaussian distributions, respectively, with a mean vector $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$. Unless otherwise specified, all indices in the paper start from 0. Moreover, we summarize the notation used in this paper in Table I.
TABLE I: Notation and definition.
Notation Definition
$N_{R}$ Number of RIS elements
$N$ Number of subcarriers
$L$ Length of the approximate PN vector
$d$ Element spacing of the RIS
$D$ Aperture of the RIS
$f_{c}$ Carrier frequency of the system
$B$ Bandwidth of the system
$\lambda$ Wavelength of the system
$c$ Speed of light
$P$ Transmit power at the AN
$\Delta f$ CFO
$\phi$ Normalized CFO
$U$ Position sampling amount
$\sigma^{2}$ Variance of the noise at the UE
$\sigma_{\Delta}^{2}$ Innovation variance for PN
$\bm{p}_{a}$ Coordinate of the AN
$\bm{p}_{r}$ Coordinate of the reference point of the RIS
$\bm{p}_{u}$ Coordinate of the UE position
$\bm{\theta}$ PN vector
$\bm{\eta}$ Approximate PN vector
$\bm{y}$ Received signal at the UE
$\bm{\mu}$ Noise-free received signal at UE
$\bm{h}$ AN-RIS-UE frequency-domain channel
$\bm{w}$ RIS phase shift vector
$\mathbf{s}$ Transmitted symbol sequence at the AN
$\bm{v}$ Additive noise at UE
$\bm{\lambda}$ Parameter vector to be estimated
$\bm{\alpha}$ Transformed parameter vector
$\mathbf{B}$ Bayesian information matrix
$\bm{\Lambda}_{\phi}$ CFO matrix
$\bm{\Lambda}_{\bm{\theta}}$ PN matrix
$\mathbf{\Pi}$ Transition matrix for the PN vector
$\mathbf{\Xi}$ Transition matrix for the HCRLB
@@ II System Model
The considered RIS-assisted OFDM downlink wireless localization system is composed of an anchor node (AN), a user equipment (UE), and an $N_{R}$-element RIS, as illustrated in Fig. 1. We make the assumption that the LoS path between the AN and UE is obstructed, and the channel state information (CSI) remains unchanged during localization, as in [21, 27, 30]. The positions of the AN and each RIS element, denoted as $\bm{p}_{a}$ and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, are known, where $r$ is the index of a RIS element. The position of UE, denoted as $\bm{p}_{u}=[x_{u},y_{u},z_{u}]^{\top}$, is to be estimated.
The RIS elements are organized in a uniform planar array (UPA). We take the bottom-left corner of the RIS as the reference point and place it at the coordinate origin. The RIS is placed on the $yz$-plane, and both edges on the coordinate axes are located in the positive half-axis. Hence, the coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times((r-1)\mod N_{R}),d\times\lceil\frac{r}{\sqrt{N_{R}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of the RIS, and is typically half of a wavelength. To satisfy the conditions of the near-field setup, both the AN-RIS distance, as well as the RIS-UE distance, should lie within the Fresnel region [26], i.e., the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{N_{R}}-1)$ represents the aperture of the RIS and $\lambda$ denotes the wavelength.
Additionally, we adopt a polar coordinate system for estimating the user’s position, i.e., using $d_{Ru}$, $\phi_{az}$ and $\phi_{el}$, to express the user’s position, indicating the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively. The relationships between these three parameters and $x_{u}$, $y_{u}$ and $z_{u}$ are given by
$d_{Ru}$ $=\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}};$ (1a)
$\phi_{az}$ $=\arctan\frac{y_{u}}{x_{u}};$ (1b)
@@
An OFDM symbol with a center carrier frequency of $f_{c}$, a signal bandwidth of $B$ and $N$ subcarriers is considered in this paper. Upon sampling at a rate of $T_{s}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at the UE, the received time-domain signal $\bm{y}$ can be written as
$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{LoS}+\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{NLoS}+\tilde{\bm{v}},$ (3)
@@
  • $P$ is the transmit power;
  • $\bm{\Lambda}_{\phi}$ is the CFO matrix resulting from a frequency mismatch between the transmitter and receiver. $\bm{\Lambda}_{\phi}$ is a diagonal matrix with $\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n\Delta f}{NB}}=e^{j2\pi\frac{n}{N}\phi}$ being the $(n,n)$-th element along its main diagonal, where $\Delta f$ is the CFO, and $\phi=\frac{\Delta f}{B}$ is called the normalized CFO;
  • $\bm{\Lambda}_{\bm{\theta}}\triangleq{\rm Diag}\left[e^{j\bm{\theta}[0]},e^{j\bm{\theta}[1]},\cdots,e^{j\bm{\theta}[N-1]}\right]$ is the PN matrix. Drawing upon the attributes of the PN observed in practical oscillators, the PN vector $\bm{\theta}$ can be modeled as a Wiener process [32], i.e.,
    $\bm{\theta}[n]=\bm{\theta}[n-1]+\bm{\Delta}[n],$ (4)
    where $\bm{\Delta}[n]$ is a Gaussian variable, i.e., $\bm{\Delta}[n]\sim\mathcal{N}(0,\sigma_{\Delta}^{2})$. As in [33], we assume that $\bm{\theta}[-1]=0$. Then, the PN vector $\bm{\theta}\triangleq\left[\bm{\theta}[0],\bm{\theta}[1],\cdots,\bm{\theta}[N-1]\right]^{\top}$ yields $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, where $\mathbf{\Psi}$ is given by
    $\mathbf{\Psi}=\sigma_{\Delta}^{2}\begin{bmatrix}1&1&1&\cdots&1&1\\ 1&2&2&\cdots&2&2\\ 1&2&3&3&\cdots&3\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&2&3&\cdots&N-1&N\end{bmatrix};$ (5)
  • $\mathbf{S}\triangleq\sqrt{N}\mathbf{F}^{\rm H}{\rm Diag}(\bm{s})$, where $\mathbf{F}^{\rm H}$ is the inverse fast Fourier transform (FFT) matrix and $\bm{s}$ is the frequency-domain transmitted symbol sequence at the AN;
  • $\bm{h}_{LoS}$ and $\bm{h}_{NLoS}$ represent the LoS and non-LoS parts of the frequency-domain channel, respectively;
  • $\tilde{\bm{v}}$ is the additive Gaussian noise at the UE, and follows the distribution $\tilde{\bm{v}}\sim\mathcal{CN}(\bm{0}_{N\times 1},\sigma_{AWGN}^{2}\mathbf{I}_{N})$.
Assume that $\bm{h}_{NLoS}$ follows a known Gaussian distribution $\bm{h}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{I}_{N})$. Thus, $\bm{\mu}_{NLoS}\triangleq\mathbf{X}\bm{h}_{NLoS}$ also follows a known Gaussian distribution $\bm{\mu}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{X}\mathbf{X}^{\rm H})$ with $\mathbf{X}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}$. Combine $\bm{\mu}_{NLoS}$ with the additive Gaussian noise $\tilde{\bm{v}}$. For conciseness, denote $\bm{h}_{LoS}$ with $\bm{h}$. The received signal is rewritten as